Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day04


Crew 102 Log Book for March 29, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: Suboptimal (headcold, stomachaches)

Time Departed/Returned from EVA: 2 EVAs 14:20 - 16:15, 17:50 - 20:00

Brief Narrative of Field Mission Results:

Collected rock samples to check for cyanobacteria existence

Repairing of the Radiotelescope Antenna

EVA Data/Interpretations:

Rock samples analysis ongoing

GPS track data being downloaded for offline analysis with dedicated software applications.

Engineering/Hab Maintenance:

Water Heater leaking problem recurring this morning.

We are currently try to save propane (low level, 13%).

Plans for Tomorrow:

Second long EVA aimed at the collection of rock samples for the cyanobacteria study.

Finalize Radiotelescope restoring via amplifier checking/calibration

Download of EVA missions GPS data (via the borrowed RS-232 - USB adapter) and their analysis on PC.

Attempt to bring back gray water system in operation based on the procedure from Mission Support Chief Engineer

Continue on current science projects (psychological, physiological, Hub Monitoring and Alarming, Hub Structural Corrosion)

Support Requested:

Waiting for propane refill.

From Observatory Team to bring radiotelescope back in operation

Miscellaneous:


Commander's Report

Franco Carbognani Reporting

We spent part of the morning in making an extra cleanup of the upper and lower floors of the Hab in order to bring dust contamination to an acceptable level. (In this respect we ask for an increase of our water consumption quota...)

After some maintenance on the ATVs (tires and oil) we started our first "long range" EVA.

Everything went smoothly but, in order to improve safety, I have decided that any other EVA with ATVs will be done using the available motorcycle helmets.

Looking carefully at them, the current EVA suite helmets may represent more a danger than

a protection in the (hopefully very unlikely) case of accident.

The EVA was productive with the collection of some cyanobacteria samples.

Even more productive was the EVA dedicated to the repair of the  Radiotelescope antenna.

Yuval was very effective at that, even with EVA suite impediments, and we are currently start looking at the signal coming to the receiver.

Mood of the crew is going lower, we have one member with an annoying headcold, two with light stomachake most probably as side effect of the "Martian" diet.

We are boiling and filtering our drinking water but still we could see inside some white particles contamination.

Lara will make a microscope analysis.



Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

    Kitty (Diesel Generator) (Used? Yes/No): Yes

    Honey (Gas Generator) (Used? Yes/No): No

    Battery Charge Level: 13.1 Volts (Measured with Voltmeter)

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

    Diesel (Full/0.75/0.5/0.25/Empty): 0.70

    Propane Tank (Full/0.75/0.5/0.25/Empty): 13%

    Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75

    Notes/Comments:

    On Mission Support Chief Engineer indications, both GreenHab termostats have been set to 50 degree F° last night in order to save propane


Oil:   

    Kitty Oil Quantity (# of Quarts in storage):

    Honey Oil Quantity (# of Quarts in storage):

    ATV Oil Quantity (# of Quarts in storage): all 4/4, fully refilled Opportunity

    Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

    Outside Potable Water Tank Level (inches from bottom): 19

    Trailer Potable Water Tank Level (inches from bottom): 32

    Hab Potable Water Tank Level (inches from bottom): 23

    Potable Water Meter Reading: 44123,4

    In to GreenHab Meter Reading:

    Notes/Comments:

Water heater leaking problem: this morning we had again some little water leaking after 15 minutes of heater on. Stopped almost immediately as soon heater switched off and warm water used.


GreenHab:

    Condition of plants in Tank 1:

    Condition of plants in Tank 2:

    Condition of plants in Tank 3:

    Greenhab Notes/Comments:    

Gray water system not functional at the moment.

Received procedure from Mission Support Chief Engineer for bringing it back in operation.

This will be performed tomorrow. 


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:

Peter Detterline and the Observatory Team has told us that the telescope need major maintenance and will be sent back to the manufacturer. We have prepared the boxes for packaging it as requested by Peter.


Transportation:


    Hab Car (NOT RENTAL)

        Used (yes/no): no

        Oil Checked (yes/no):yes

        Tire Status: good

        Odometer at end of day: 123761.4

        Notes/Comments on Hab Car:


    ATV 1 (Opportunity):

        Used: y

        Oil Checked: y (refilled)

        Fuel Consumed: 1/2 liter

        Tire Status: inflated

        Comments:


    ATV 2 (Spirit):

        Used: no

        Oil Checked: y

        Fuel Consumed: none

        Tire Status: inflated

        Comments:


    ATV 3 (Viking I):

        Used: y

        Oil Checked: y

        Fuel Consumed: 1/2 liter

        Tire Status: good

        Comments:


    ATV 4 (Viking II):

        Used: y

        Oil Checked: y

        Fuel Consumed: 1/2 liter

        Tire Status: good

        Comments:


Heating and Ventilation:

    Thermostat setting upstairs (Farenheit): 67

    Actual temperature upstairs (Farenheit): 73

    Thermostat setting downstairs (Farenheit):65

    Actual temperature downstairs (Farenheit): 62


Computers/Networking Infrastructure:

    Notes/comments on internet/computers?:


General Engineering Notes/Comments:

The setup of the alarming system (F. Carbognani science project) is continuing.

A web server is now running on the Linux notebook and corresponding web pages reporting in real time: Temperature, CO2 concentration and Humidity in the Hab are accessible on all local computers.



Health and Safety Officer's Report

Dr. John E. Deaton Reporting

No significant issues to report at this time. Health and Safety Officer has contracted a head cold and cough, but it seems to be manageable and not increasing in severity. One dose of Theraflu was administered last evening prior to retiring.

Would advise modifications to helmets worn during EVAs. Protrusions of a type of bolt or screw are evident on outside of helmets which could prove to be hazardous if personnel stumble and/or incur an accident during an EVA using the ATVs.


Science Report (Space Florida Outreach Experiment - Geology)

Kavya Manyapu Reporting

OBJECTIVE

The objective of this project was to inspect if the sublimation of dry ice can be insulated in silica. The scientific relevance of this investigation is to further the exploration of life by studying the properties of frozen carbon dioxide and the impact on Martian soil.

PROCEDURE

1.      A block of dry Ice measuring  5in x 3in X  1in (total volume of 15 in3 and total surface area of  46 in2 was used for this experiment

2.      The terrain outside the habitat was inspected  to find a convenient spot to bury the dry ice that is free of rocks

3.      For the first trial a hole of 6 inches was dug into the soil and the dry ice was placed in the hole. The hole was covered with 1 in of sand and the sublimation rate was recorded for 3 30s intervals

4.      A second trial with 3 in sand to cover the dry ice was used and the same process was followed

5.      Eventually, the hole was covered with 10inches and the sublimation rate of dry ice was inspected

6.      The soil was extracted to see if the dry ice was fully sublimated after a period of 3 hours

RESULTS

The first trial with 1 in sand did not insulate the dry ice sublimation to a large extent as fumes of dry ice were visible from the surface. But when the dry ice was covered with 10 inches of sand the dry ice vapors were no longer seen. This shows that silica has an ability to insulate the sublimation process.  The soil was extracted sublimation process. It is suspected that the soil will have the sublimation to a large extent as fumes of dry ice were visible from the surface. But when the dry ice was covered with 10 inches of sand the dry ice vapors were no longer seen. This shows that silica has an to see if the dry ice was still present. After a period of 3 hours a small amount of dry ice (2 in x 1 in x 0.2 in) was still present. This proves that silica has the capabilities of slowing down the dry ice capability to insulate dry ice, but the depths at which this can be completely achieved are still to be determined experimentally.

However, the complete insulation of dry ice by silica was not completely tested as this requires digging a hole as deep at 1-2 meters. Because of operational constraints this could not be accomplished.

LIMITATIONS

•       Because of the visibility constraints due to the space suit helmets, the dry ice vapor fumes on the surface of the soil were very hard to detect. This might have led to some bias in the data collected. The soil was covered from the top with a plexiglass container in order to be able to see the fumes (Figure 1).

•       The location had rain showers the night before which saturated the soil. The saturation of the soil the next day was not known and this might have also contributed to the insulation of the dry ice vapors.

Soil saturation as a result of the water (rain) may have increased the insulation capability of the soil.  Hence, the experiment should be repeated when the soil is completely dry. However, because of the unavailability of dry ice at analogue Mars, this experiment could not be repeated.


Science Report (Space Florida Outreach Experiment - Biology)

Lara Alaska Reporting

SURVIVAL OF TARDIGRADES TO EXTREMELY COLD TEMPERATURES

Abstract:

Tardigrades, also known as Water Bears, are microscopic animals that display broad resistance to a complex interplay of stressful conditions. They have been found able to tolerate extreme temperatures, high UV light exposure as well as low water availability. Under dry conditions they are able to enter a state of suspended animation termed cryptobiosis where the animal’s body water percentage will decrease to less than 1%. Many species of the Tardigrade phylum are characterized by an increased resistance to very low temperatures when they are in a dehydrated state. Their hardy nature makes them suitable organisms to test resistance of life in space studies and has made them deserve the name of polyextremophiles . Our objective was to test whether viable Tardigrades survive exposure to extremely low temperatures for different amounts of time.


Procedure:

In our study samples of Tardigrades were subjected to extremely cold temperatures provided by direct contact with dry ice. Each sample (a few microliters of the main culture) put on a slide was first checked for viability under the microscope. Tardigrades in each sample displayed classical morphology with a short barrel-shaped body and well visible legs and claws. The population seemed to be heterogeneous as the organisms presented a wide range of sizes. They were very mobile and were seen in a feeding mode within masses of algae. A large number of ciliates feeding on algae was also visible.

Each sample was subjected to direct contact on top of dry ice placed inside a cooler for different amounts of time (3 seconds, 10 seconds,15 seconds, 30 seconds, 1 minute). The samples were then retrieved and viability was then checked under the microscope every 10 minutes up to 1 hour after exposure.


Results:

The presented study had different outcomes on viability of Tardigrades depending upon the exposure time and on the proximity of the slide to the heterogeneous surface of the dry ice:

-Sample 1 (1 minute exposure) showed how all organisms, both Tardigrades and ciliates, died following to total freezing of the slide. No viable organism was seen at any time after the exposure.

- Sample 2 to 4 (10, 15, 30 seconds) showed no viability of organisms that had been in direct contact with the dry ice (the slide was frozen on those areas), while those organisms that were in areas of the slide that were in proximity, but not in direct contact with the dry ice, displayed partial mobility that increased after letting them recover for 1 hour.

-Sample 5 (3 seconds) showed that no organism had suffer from the cold shock and the time was not enough to freeze the slide.


Discussion and future work:

The experiment carried out to test Tardigrades resistance to cold temperature showed that these animals do not survive to direct contact to dry ice. This result is probably due to the fact that placing the slide on direct contact to dry ice caused them to freeze nearly instantaneously leaving no time for the organism to adjust to the new temperature. Many Tardigrade species, if cooled gradually, form a cold-resistant tun which allows the animal to survive being frozen.

Those slides that displayed partial viability in their organisms were probably due to the fact that certain areas of the slide had not frozen due to the irregular surface of the dry ice available and viable organisms were accordingly seen in non-frozen areas.

Future experiments should therefore allow a gradual decrease of temperature to which the organisms are subjected to test cold resistance. A comparative quantitative study could also be performed to test whether samples previously dehydrated show a higher resistance to cold temperatures compared to hydrated organisms.


Science Report (Human Factors/Psychology)

Dr. John E. Deaton Reporting

Human performance tests focusing primarily on cognitive abilities (sleepiness, math processing, pattern matching, mood, and logical processes) continue to be administered to crew members each day. These tests will be repeated every other day until the completion of the mission.

The NASA TLX Task Load Index was administered to three crew members prior to an EVA and subsequent to the completion of an EVA to evaluate workload levels sustained during the 2 hour mission that occurred at approximately 1430 today.  A second EVA to repair the radio antenna was another opportunity today to administer the NASA TLX to measure workload levels resulting in the repair of the radio antenna.


EVA 3 Report

Yuval Brodsky Reporting

Time: 1420 - 1615

Crew Members: Y. Brodsky (EVA commander), F. Carbognani, L. Vimercati

Site Location: Fossil rich site ~1.0 mi bearing 075 from MDRS, fossil rich site ~0.8 mi bearing 345 from MDRS

MDRS veteran L. Vimercati introduced other EVA participants to fossil-rich locations.

Transit Mode: ATVs

Primary objective: Collection of cyanobacteria-rich rock samples.

Secondary objective: Exploration of fossil-rich environments.

Results: Crew gathered ~350 grams of rock containing cyanobacteria, of which <5% (mass) is cyanobacteria. Further EVAs will be required to gather additional samples. All EVA participants familiarized with fossil-rich environments surrounding MDRS. Crew had uneventful encounter with rattlesnake while on EVA.

Lessons Learned: EVA suit helmets provide inadequate visibility for safe ATV operation. ATVs should be operated using motorcycle helmets, and EVA helmets donned upon arrival at EVA site.


EVA 4 Report

Yuval Brodsky Reporting

Time: 1750 - 2000

Crew Members: Y. Brodsky (EVA commander), A. Kapoglou, K. Manyapu

Site Location: Radiotelescope antennae 20 meters East of Hab

Transit Mode: Foot

Objectives: Inspection of damaged radiotelescope antennae, repair of damaged antennae.

Results: Both dipole antennae raised and suspended between their PVC pipes to height of ~10m. All four PVC pipes supporting antennae righted to vertical position, secured with guy-lines. Antennae suspected to be in an operational configuration, radiotelescope to be tested in coming 24 hrs. Second EVA may be required to adjust antennae positions. Visor fogging introduced challenges during this EVA.

Lessons Learned: Very helpful to have HabCom with EVA procedures manual and checklists. 


Chef's Report

Kavya Manyapu Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) were prepared to mark the celebration?

Just another day

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Leonardo da fettuccine

Mushroom pilaf with Vegetables

Western Tamale Pie

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

1 servings Leonardo de fettuccine

2 servings Western Tamale

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

5) Lessons learned (if any)

The crew did not enjoy the leonardo da fettuccine at all.

6) Comments/questions for Kim and Jean, the food study investigators

The leonardo da fettuccine was not at all well received by ALL the crew. It is suggested that this be removed from the food items. Some of the crew started having stomach aches and they believe that it is because of the type of food in the habitat.


Journalist Report

Kavya Manyapu Reporting

Perhaps the most exciting thing that happened today for the crew was spotting a Martian alien that looked like a rattle snake during their EVA.  Dust has been a major problem here on Mars. Hence the day started out with cleaning and mopping the hab completely as much as the crew could to sanitize the living area. This activity was then followed by lunch.  Today was a non cooking day, meaning that the food is already pre-prepared and redyhrated and needs to be boiled in order to consume. Most of the crew did not like the pre-prepared food , although the dish was Leonardo da Fettuccine, as it did not have the slightest taste of what it claims!

The first EVA of the day was commanded by Y.Brodsky, accompanied by Crew Biologist L.Vimericati and Crew Commander F.Carbognani. The cap-com for this EVA was the crew Engineer/Physiologist K.Manyapu. The EVA involved collecting cyanobacteria samples and also to look for dianosaur fossils at 3.6 miles away from the habitat. The crew traversed the Martian area using Extraterrestrial Vehicles. The crew came back safe with some samples. EVA 1 lasted 2 hours and was physically demanding . The  second EVA was commanded by Y.brodsky accompanied by K.Manyapu and A.Kapoglou. The EVA involved the repairing of the radio telescope and also taking further samples of the habitat structure for corrosion testing. The crew continues to do exciting things and contributing to science by taking the up the challenge of living on an isolated planet and working their way through. Keep reading these reports for more exciting things to come.

Until then a big bye from Mars!